OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4. y = x + 2 y = -x - 2 y = -x + 4 y = x - 4

OpenStudy (triciaal):

parallel lines have the same slope given slope, m is -1 you have the point (-3, 1) let this be (x1, y1) point slope formula (y - y1) = m(x- x1) plug your numbers and simplify put in the form y = mx + b to match with ans choices